# DecoderTracker: Decoder-Only Method for Multiple-Object Tracking

### v4

@@ 3.3 Loss and Training Methods
The third stage of training involves the utilization of TALA during the training phase, while simultaneously deactivating the TBSP. Experimental evidence indicates that if the training epochs during this stage are sufficient, TBSP can also be disabled during the inference phase without significantly compromising performance. This finding underscores the robustness of the model, suggesting that once adequately trained in the third stage, TBSP can be turned off during inference with negligible impact on overall performance.
@@ 4.1 Datasets and Metrics
To validate DecoderTracker’s performance, we evaluate the algorithm on three demanding datasets: DanceTrack[32], MOT17[33], and KITTI[34].
DanceTrack is a vast dataset designed for human tracking, showcasing scenarios with occlusion, frequent crossovers, uniform appearances, and varied body gestures. Comprising 100 videos featuring diverse dance styles, it underscores the significance of motion analysis in multi-object tracking.
@@
Evaluation Metrics: We assessed our method using widely recognized MOT evaluation metrics. The primary metrics employed for evaluation included HOTA, AssA, DetA ,IDF1 and MOTA[35, 36]. These metrics collectively provided a comprehensive and robust evaluation of our model’s tracking performance. HOTA assessed the tracking accuracy based on the spatial and temporal overlap between predicted and ground truth tracks. AssA measured the quality of associations between objects, while IDF1 quantified the accuracy of identity tracking. MOTA offered an overall evaluation of tracking accuracy, considering false positives, false negatives, and identity switches. These metrics formed the basis for a thorough evaluation of our model’s tracking capabilities.
@@ 4.2 Implementation Details
The rationale for adopting DecoderTracker originates from our observation of the high training costs and suboptimal inference speed associated with MOTR. Thus, our aim is to propose a new baseline model surpassing MOTR’s performance. To ensure a fair comparison, we exclusively employ methods enhancing tracking performance, while omitting improvement techniques utilized in MeMOTR [10], MOTRv2 [18], MOTRv3 [9], and CO-MOT [8].
The majority of experiments were conducted using PyTorch on a single NVIDIA GeForce RTX 4090 GPU. For equitable comparison with MOTR, training on the MOT17 dataset was performed on a single NVIDIA GeForce RTX 2080ti. Additionally, to better contrast with the MOTR series in terms of speed, inference speed tests were carried out on a single NVIDIA V100 GPU. One more thing, he image size for the input network is $640\times 640$.
During the initial training, batch sizes were set to 32 on the 4090 and 16 on the 2080ti. Subsequent stages utilized a uniform batch size of 1. For MOT17, like MOTR, we integrated extra data from the CrowdedHuman [37], employing random displacement to generate video segments with pseudo-trajectories. Training comprised 120 epochs in the first stage, followed by 30 epochs in the second. The third stage extended to 20 epochs with TBSP, or 55 without. For Dancetrack, ensuring parity without extra data like MOTRv2, the first stage was 15 epochs, the second was 9, and the third, if TBSP was used, was 9 epochs, or 18 without TBSP. For KITTI, initial training involved 80 epochs, followed by 25 in the second stage. The third stage consisted of 30 epochs with TBSP or 70 without. Additionally, approximately 5k images from BDD100k [38] were selected for additional training.
4.3 Performance Evaluation and Comparison
In Table 1, a comprehensive comparison of training time and speed was conducted between DecoderTracker and the MOTR series. It is crucial to note that the datasets used for comparing training time and inference speed exhibit variations. This disparity primarily arises from the act that the publicly disclosed training times and inference speeds of the MOTR series are based on different datasets.
@@ 4.3 Performance Evaluation and Comparison
Table 1: Comparison of training time (including all previous stages) and speed between DecoderTracker and MOTR series on a single V100 GPU. The data for the MOTR series is sourced from [3], [9] or their open source address.
DecoderTracker demonstrates a substantial advantage over the MOTR series, particularly in terms of training time and speed. The primary reason behind this efficiency gain lies in the architectural simplicity of DecoderTracker. Unlike MOTR, DecoderTracker adopts a streamlined structure that eliminates the need for an encoder, resulting in a more computationally frugal design. This reduction in complexity enables DecoderTracker to achieve training completion within 68.7 hours, utilizing only 1 Nvidia GeForce 2080ti GPU. In contrast, MOTR relies on 8 Nvidia GeForce 2080ti GPUs and takes about 96 hours for training. The absence of an encoder in DecoderTracker contributes to its swift training process, emphasizing the model’s efficiency in resource utilization.
Table 2: Performance comparison between DecoderTracker and existing methods on the Dancetrack[32] benchmark. Numbers are highlighted in bold when they correspond to either DecoderTracker or the superior metric from MOTR.
Methods HOTA IDF1 DetA AssA MOTA
@@
ByteTrack[14] 47.3 52.5 71.6 31.3 89.5
GTR[39] 48.0 50.3 72.5 31.9 84.7
DST-track[40] 51.9 51.0 82.3 34.6 84.9
OC-SORT[41] 54.1 54.2 80.4 38.0 89.4
@@
MeMOTR[10] 63.4 65.5 77.0 52.3 85.4
MOTR[3] 54.2 51.5 73.5 40.2 87.6
DecoderTracker(Ours) 54.2 54.2 78.7 37.4 87.8
On the Dancetrack test benchmark, DecoderTracker(TBSP) achieves a speed of 18.5 frames per second (FPS) using a single V100, while the MOTR series attains speeds of 9.5 FPS (MOTR), 6.9 FPS (MOTRv2), and 10.6 FPS (MOTRv3). Particularly noteworthy is the remarkable 19.6 FPS achieved by DecoderTracker, emphasizing its outstanding performance in inference speed. Table 2, Table 3, and Table 4 exhibit the final results of end-to-end DecoderTracker on the Dancetrack, MOT17, and KITTI benchmark, respectively. In contrast to MOTR, DecoderTracker demonstrates competitive performance across multiple metrics on both datasets. It is important to highlight that the performance of MeMOTR is provided in Tables 2 and 3. The inclusion of MeMOTR’s performance aims to underscore the considerable potential for advancement and enhancement in both MOTR and our DecoderTracker. MeMOTR[10] integrates additional modules and performance-enhancing techniques, along with supplementary datasets during training. Consequently, comparing its tracking performance directly with that of DecoderTracker and MOTR may not be entirely fair.
Table 3: Performance comparison between DecoderTracker and existing methods on the MOT17[33] set.
method HOTA IDF1 DetA AssA MOTA
@@
FairMOT[42] 59.3 72.3 60.9 58.0 73.7
ByteTrack[14] 62.8 77.2 63.8 62.2 78.9
StrongSORT[39] 63.5 78.3 63.6 63.7 78.5
UTM[43] 64.0 78.7 65.9 62.5 81.8
OC-SORT[41] 63.2 77.5 63.2 63.2 78.0
Deep OC-SORT[6] 64.9 80.6 64.1 65.9 79.4
@@
MeMOTR[10] 58.8 69.0 59.6 58.4 72.5
MOTR[3] 57.2 59.4 58.9 55.8 71.9
DecoderTracker(Ours) 55.7 70.5 55.7 56.0 66.5
For Dancetrack (Table 2), DecoderTracker achieves comparable results to MOTR, with same scores in HOTA but outperforming MOTR in DetA. Specifically, DecoderTracker achieves an impressive MOTA, indicating its overall superior performance in multi-object tracking tasks.

### v5

@@ 3.1.5 Loss and Training Methods
The third stage of training involves the utilization of TALA during the training phase, while simultaneously deactivating the TBSP. Experimental evidence indicates that if the training epochs during this stage are sufficient, TBSP can also be disabled during the inference phase without significantly compromising performance. This finding underscores the robustness of the method, suggesting that once adequately trained in the third stage, TBSP can be turned off during inference with negligible impact on overall performance.
3.2 DecoderTracker+
Following the improvements in previous sections that led to the proposal of DecoderTracker, we observed suboptimal enhancements in inference speed, failing to meet anticipated performance benchmarks. Through latency profiling of individual modules, we identified the decoder as the primary bottleneck limiting acceleration. Subsequent in-depth analyses of GPU memory utilization and query quantity revealed that the core issue likely stems from dynamic data structures. Compared to static data configurations, dynamic data introduces inference challenges such as inefficient memory allocation scheduling, with empirical evidence provided in Figure 6.
To address this, we present DecoderTracker+ in this chapter. By implementing a Fixed-Size Query Memory(FSQM) mechanism that maintains a constant number of tracking queries, our approach enables deeper optimization compatibility with modern deep learning frameworks ,e.g., PyTorch, TensorFlow. Furthermore, the deterministic data structure facilitates straightforward conversion to TensorRT, paving the way for accelerated deployment of end-to-end MOT methods on edge devices.
DecoderTracker+ maintains the core architectural framework and operational workflow of its predecessor, as systematically illustrated in Figure 2, while introducing strategic innovations primarily in two key dimensions: Fixed-Size Query Memory and Attention Computation Optimization. Subsequent subsections provide detailed technical expositions of these advancements.
Figure 6: The two figures illustrate the changes in Decoder memory occupancy and the latency of various components for MOTR, its base model Deformable DETR, our DecoderTracker, and DecoderTracker+ on the MOT17-14 sequence. In the right figure, the backbone of our methods is represented by FENet, and since they do not have an Encoder, their latency on the Encoder is 0. The memory changes of DecoderTracker are basically consistent with those of MOTR, so they are not displayed here. It can be clearly seen from the figures that dynamic data leads to unstable memory occupancy of the Decoder and abnormally high latency.
3.2.1 Fixed-Size Query Memory
FSQM maintains fixed-dimensional queries $q_{tr}^{\mathcal{M}}\in\mathbb{R}^{N\times d}$, where $N$ denotes the preset maximum number of tracking targets and $d$ represents the feature dimension. Compared to dynamic query mechanisms, this approach introduces:
$\hat{\omega}_{tr}^{\mathcal{M}}=(\underbrace{q_{tr}^{\mathcal{M}}}_{\text{static queries}},\underbrace{C_{\text{tr}}\in\mathbb{R}^{N}}_{\text{confidencevector}},\underbrace{\mathcal{ID}}_{\text{global ID pool}},\underbrace{BBox^{\mathcal{M}}}_{\text{bounding box}})$ (5)
to model memory for the continuous tracking process. During initialization, set $q_{ftr}^{\mathcal{M}}=\mathbf{0}$ and $C_{\text{tr}}^{(0)}=\mathbf{0}$ for all queries, where all queries are initialized to zero to avoid interference with networks that do not rely on attention mechanisms. Additionally, all inactive queries are also initialized to zero. Moreover, a global ID pool, denoted as $\mathcal{ID}$, is introduced. Each active tracking query is assigned a unique ID that accompanies the query as it moves across video frames. When a query is no longer associated with a tracked object, its corresponding ID is deleted and will not be reused. A new ID is assigned to new active queries. By the way, all inactive queries are assigned an ID of -1 and their bounding boxes are initialized with zero values.
During each frame processing, iterative updates are performed as follows:
  1. 1.
    New query injection: For detection queries assignment $\{\hat{\omega}_{\text{det}}^{(i)}\}$ with detection confidence $c_{i}>\tau_{\text{in}}$, each query is directly assigned the first available ID from the global ID pool (which initially contains the value -1 for new queries). The corresponding query in the memory pool is updated using a momentum update rule:
    $j=\mathop{\arg\min}\limits_{k}(\mathcal{ID}[k]=-1),$ (6)
    $\hat{\omega}_{tr}^{\mathcal{M}}{(t)}[j]=\hat{\omega}_{\text{det}}^{(i)}$
    After updating the query, its corresponding ID is assigned from the global pool.
  2. 2.
    Query removal: For the existing queries in the memory pool, confidence is monitored over time. If a query’s confidence falls below a threshold $\tau_{\text{out}}$ for three consecutive frames, it is considered for removal. Upon removal, the query is reset to zero, and its corresponding ID is set to -1. Specifically:
    $\text{If }C_{\text{tr}}^{(t-3)}[n]<\tau_{\text{out}},C_{\text{tr}}^{(t-2)}[n]<\tau_{\text{out}},C_{\text{tr}}^{(t-1)}[n]<\tau_{\text{out}},$ (7)
    then the query is removed by setting:
    $q_{tr}^{\mathcal{M}}[n]=\mathbf{0},\quad\mathcal{ID}[n]=-1.$ (8)
    where $S(\cdot)$ represents the prediction score function. The decay ensures that the relevance of older queries decreases over time, while new, active queries retain higher confidence scores.
It is also important to emphasize that, aside from this, the processing of queries in FSQM is consistent with that in DecoderTracker. The specific process will not be repeated here, but can be referred to in the previous sections.
Algorithm 2 FSQM Online Update Procedure
0:  New query set $\hat{\omega}_{\text{det}}=\{q_{i},c_{i},BBox_{i}\}_{i=1}^{M}$, memory pool $\hat{\omega}_{tr}^{\mathcal{M}}=(q_{\text{tr}},C_{\text{tr}},\mathcal{ID},BBox_{tr})$
0:  Updated memory pool $\mathcal{M}^{\prime}$
1:  $\hat{\omega}_{\text{valid}}\leftarrow\{\hat{\omega}_{i}|c_{i}>\tau_{\text{in}}\}$Filter valid new queries}
2:  if $\hat{\omega}_{\text{valid}}\neq\emptyset$ then
3:     for $q_{j}\in\mathcal{\hat{\omega}}_{\text{valid}}$ do
4:        Assign the first available ID from the global ID pool $\mathcal{ID}$ to $q_{j}$
5:        $k\leftarrow\mathop{\arg\min}\limits_{n}\mathcal{ID}[n]=-1$Find the first inactive query}
6:        $q_{\text{tr}}[k]\leftarrow q_{j},BBox_{\text{tr}}[k]\leftarrow BBox_{j}$Momentum update}
7:        $C_{\text{tr}}[k]\leftarrow c_{i}$Reset confidence to 1.0}
8:     end for
9:  end if
10:  for $n\in\{1,...,N\}$ do
11:     if $C_{\text{tr}}[n]<\tau_{\text{out}}$ for 3 consecutive frames then
12:        $q_{\text{tr}}[n]\leftarrow\mathbf{0},BBox_{\text{tr}}[k]\leftarrow\mathbf{0}$Set query and BBox to zero (inactive)}
13:        $\mathcal{ID}[n]\leftarrow-1$Reset ID to -1}
14:     end if
15:  end for
16:  return $\mathcal{M}^{\prime}=(Q_{\text{tr}},C_{\text{tr}},\mathcal{ID})$
3.2.2 Attention Mechanism Optimization
Since some queries in FSQM are empty, we need to explore the network structure to ensure that these empty queries do not affect the final results. The network processing these queries consists only of linear layers, normalization layers, dropout, activation function, cross-attention with image features, and finally, self-attention. Our first step is to analyze how empty queries may impact the output of these networks.
Firstly, we will analyze some unaffected networks. These network architectures inherently avoid commingling active and inactive queries in their computational processes. Therefore, during inference, there is no risk of interference arising from their interactions. Our subsequent analysis need only focus on verifying whether these networks produce gradients during the training.
For linear layers, zero-valued queries produce the following output:
$W\mathbf{0}+b=b$ (9)
Here, $W$ is the weight matrix, $\mathbf{0}$ is the zero-valued query, and $b$ is the bias term. Since the zero-valued query contributes nothing, only the bias remains, which can be neutralized through layer normalization. After applying layer normalization, the output is given by:
$\hat{y}=\frac{y-\mu}{\sigma}\cdot\gamma+\beta$ (10)
where $y=b$, $\mu$ and $\sigma$ are the mean and standard deviation of the output across the batch, and $\gamma$ and $\beta$ are learnable parameters. Since the bias term $b$ is small in practice, empirical tests show that its $L_{2}$ norm $\|b\|_{2}$ is very small, with $\|b\|_{2}<0.03$ in our implementation.
For normalization layers, since the zero-valued queries do not produce any meaningful activation, they do not affect the overall normalization process. The query’s zero value is normalized along with other queries, resulting in minimal influence on the normalized output. This can be expressed as:
$q_{i}=\frac{q_{i}-\mu}{\sigma}\cdot\gamma+\beta\quad\text{for active queries }q_{i}$ (11)
where $\hat{q}_{i}$ is the normalized output for an active query $q_{i}$. For inactive queries (those initialized to zero), the effect is negligible due to their zero contribution.
For dropout, which randomly sets a fraction of the inputs to zero during training, inactive queries do not contribute to the dropout process. Inactive queries, having values of zero, remain unaffected by dropout, and the overall output for those queries is still zero. Mathematically, dropout can be expressed as:
$\text{Dropout}(\omega_{i})=\begin{cases}0&\text{if query is inactive (}q_{i}=0\text{)}\\ q_{i}\cdot d&\text{if query is active}\end{cases}$ (12)
where $d\in[0,1]$ is a random variable representing the dropout rate. Inactive queries (zero-valued) remain zero regardless of whether dropout is applied.
For activation function, since only GELU is used to process queries, we will only discuss GELU here. In the case of GELU, the output for a zero-valued query is:
$\text{GELU}(q_{i})=0.5\omega_{i}\left(1+\tanh\left(\sqrt{\frac{2}{\pi}}(q_{i}+0.044715q_{i}^{3})\right)\right)\quad\text{for }q_{i}=0\implies\text{GELU}(0)=0.$ (13)
Thus, for activation function, zero-valued queries result in zero output, ensuring that they do not influence the subsequent layers.
For cross-attention, given key matrix $K$ and dimension $d$ of Query, the attention scores become:
$\text{Attn}(Q_{\text{zero}},K)=\frac{\mathbf{0}\cdot K^{T}}{\sqrt{d}}=\mathbf{0}$ (14)
resulting in zero contribution after softmax normalization.
In summary, for linear, normalization, dropout, activation layers and cross-attention, zero-valued queries contribute nothing to the computation, ensuring that they do not interfere with the active queries during the forward and zero-valued in the forward inherently prevent gradient computation, thereby eliminating any impact on the training.
Regarding the self-attention layers, since the queries, keys, and values used in computing self-attention are all derived from the queries themselves, inactive queries inevitably introduce interference to the output. While a detailed analysis of this phenomenon is omitted here, we address this issue effectively by implementing a masking mechanism. Here we use the self attention layer in Decoder to illustrate, and the similarity in TAN will not be further elaborated.
When calculating self-attention, we first compute an attention weight matrix $A$:
$A=\text{softmax}\left(\frac{q(q^{\prime})^{T}}{\sqrt{d}}+\log(\text{Mask})\right)$ (15)
where $q\in\mathbb{R}^{2N\times d}$ is obtained by concatenating $q_{tr}^{\mathcal{M}}$ from FSQM and $q_{det}$, i.e., $q=\text{concat}(q_{tr}^{\mathcal{M}},q{det})$ in first decoder layer, while in other layers, $q$ is the output from the previous layer. $(\text{Mask}\in\mathbb{R}^{2N\times 2N}$ is the mask we need to generate. Its dimensions are the same as those of $A$, and it ensures that after applying the softmax function, the unwanted values are set to zero. By appropriately using this mask during the final self-attention output $A\times V$ where $V$ is derived from $Q$ with positional encoding, the effect of empty queries on the final result can be effectively eliminated.
Here, we extend $\mathcal{ID}$ and then generate the mask.
$\mathcal{ID}_{\text{ext}}=\text{Concat}(\mathcal{ID}_{\text{tr}},\mathbf{0}_{N})$ (16)
where, $\mathbf{0}_{N}$represents a fixed identifier for detect queries.
$\text{Mask}_{ij}=\begin{cases}0&\text{if }\mathcal{ID}_{ext}[i]=-1\lor\mathcal{ID}_{ext}[j]=-1\\ 1&\text{otherwise}\end{cases}$ (17)
By implementing this mask mechanism, we ensure that values associated with inactive queries in matrix $A$ are nullified to zero, thereby eliminating their influence on the final output. Notably, as demonstrated through our preceding analyses, the inactive query positions inherently maintain zero values across all decoder layers. Consequently, this masking mechanism uniformly applies to self-attention operations in every decoder layer without requiring layer-specific adaptations.
In terms of the mask in the self-attention layer of the TAN, the generation process and mechanism are largely consistent with those described earlier; however, the dimensions in TAN are fewer, and there is no need to expand $\mathcal{ID}$.
@@ 4.1 Datasets and Metrics
To validate DecoderTracker and DecoderTracker+’s performance, we evaluate the algorithm on three demanding datasets: DanceTrack[37], MOT17[38], and KITTI[39]. DanceTrack is a vast dataset designed for human tracking, showcasing scenarios with occlusion, frequent crossovers, uniform appearances, and varied body gestures. Comprising 100 videos featuring diverse dance styles, it underscores the significance of motion analysis in multi-object tracking.
@@
Evaluation Metrics: We assessed our method using widely recognized MOT evaluation metrics. The primary metrics employed for evaluation included HOTA, AssA, DetA ,IDF1 and MOTA[40, 41]. These metrics collectively provided a comprehensive and robust evaluation of our method’s tracking performance. HOTA assessed the tracking accuracy based on the spatial and temporal overlap between predicted and ground truth tracks. AssA measured the quality of associations between objects, while IDF1 quantified the accuracy of identity tracking. MOTA offered an overall evaluation of tracking accuracy, considering false positives, false negatives, and identity switches. These metrics formed the basis for a thorough evaluation of our method’s tracking capabilities.
@@ 4.2 Implementation Details
The rationale for adopting DecoderTracker originates from our observation of the high training costs and suboptimal inference speed associated with MOTR. Thus, our aim is to propose a new baseline method surpassing MOTR’s performance. To ensure a fair comparison, we exclusively employ methods enhancing tracking performance, while omitting improvement techniques utilized in MeMOTR [9], MOTRv2 [18], MOTRv3 [8], and CO-MOT [7].
The majority of experiments were conducted using PyTorch on a single NVIDIA GeForce RTX 4090 GPU. For equitable comparison with MOTR, training on the MOT17 dataset was performed on a single NVIDIA GeForce RTX 2080ti. Additionally, to better contrast with the MOTR series in terms of speed, inference speed tests were carried out on a single NVIDIA V100 GPU. One more thing, the image size for the input network is $640\times 640$.
@@
DecoderTracker+ 28.8
Table 1: Comparison of training time (including all previous stages) and speed between DecoderTracker and MOTR series on a single V100 GPU. The training time of DecoderTracker+ is basically the same as that of DecoderTracker, so it will not be listed separately here.
Figure 7: IDF1-MOTA-HOTA and IDF1-MOTA-FPS comparisons of state-of-the-art trackers with our proposed DecoderTracker and DecoderTracker+ on Dancetrack test sets. For the left image, the horizontal axis is MOTA, the vertical axis is IDF1, and the radius of the circle is HOTA. For the right image, the horizontal axis is FPS, the vertical axis is IDF1, and the radius of the circle is HOTA. ”*” represents additional data was used for training. Our DecoderTracker+* achieves the best IDF1 and HOTA and comparable speed performance.
Table 2: Performance comparison between DecoderTracker and existing methods on the Dancetrack[37] test set. Numbers are highlighted in bold when they correspond to either DecoderTracker or the superior metric from MOTR. * means that extra data was used during training.
Methods IDF1 HOTA AssA DetA MOTA
@@
TransTrack[15] 45.2 45.5 27.5 75.9 87.7
GTR[21] 50.3 48.0 31.9 72.5 84.7
DST-track[42] 51.0 51.9 34.6 82.3 84.9
ByteTrack[14] 52.5 47.3 31.3 71.6 89.5
OC-SORT[22] 54.2 54.6 38.0 80.4 89.4
Deep OC-SORT[5] 61.5 61.3 45.8 82.2 92.3
@@
MOTR[2] 51.5 54.2 40.2 73.5 79.7
DecoderTracker 54.2 54.2 35.7 78.7 87.8
DecoderTracker+ 54.4 54.5 35.6 79.0 88.1
DecoderTracker* 62.9 60.8 47.1 81.2 91.1
DecoderTracker+* 62.2 61.2 46.0 81.8 91.9
Table 3: Performance comparison between DecoderTracker and existing methods on the MOT17 [38] set. * means that extra data was used during training.
method IDF1 HOTA DetA AssA MOTA
@@
YOLOTracker[43] 65.1 - - - 71.9
FairMOT[44] 72.3 59.3 60.9 58.0 73.7
MSPNet[45] 74.6 - - - 71.9
ByteTrack[14] 77.2 62.8 63.8 62.2 78.9
OC-SORT[22] 77.5 63.2 63.2 63.2 78.0
StrongSORT[21] 78.3 63.5 63.6 63.7 78.5
UTM[46] 78.7 64.0 65.9 62.5 81.8
Deep OC-SORT[5] 80.6 64.9 64.1 65.9 79.4
@@
MOTR[2] 59.4 57.8 60.3 55.7 73.4
DecoderTracker 70.5 55.7 55.7 56.0 66.5
DecoderTracker+ 70.4 55.8 55.2 55.8 66.7
DecoderTracker* 77.2 61.8 62.1 60.1 76.8
DecoderTracker+* 77.5 62.2 62.5 60.7 76.7
During the initial training, batch sizes were set to 32 on the 4090 and 16 on the 2080ti. Subsequent stages utilized a uniform batch size of 1.
For MOT17, like MOTR, we integrated extra data from the CrowdedHuman [47], employing random displacement to generate video segments with pseudo-trajectories. Training comprised 120 epochs in the first stage, followed by 30 epochs in the second. The third stage extended to 20 epochs with TBSP, or 55 without. And, to address the issue of the limited scale of the MOT17 dataset, this paper employs a hybrid dataset. It integrates the MOT20 dataset and the complete CrowdHuman dataset, which has been spatially shifted, to expand data diversity. The amount of new data significantly exceeds the benchmark set (by approximately 4.3 times). To balance training bias, a stratified sampling mechanism was designed: within each epoch, samples from the MOT17 dataset are resampled twice, while samples from the supplementary sets are sampled once. This approach not only alleviates data skewness but also enhances the learning of benchmark set features, ensuring that the model maintains its original performance advantages while improving its adaptability to dense scenes.
For Dancetrack, ensuring parity without extra data like MOTRv2, the first stage was 15 epochs, the second was 9, and the third, if TBSP was used, was 9 epochs, or 18 without TBSP. To better compare with the state-of-the-art (SOTA) methods, when testing on the test set, we also used the CrowdHuman dataset as supplementary data and the validation set of Dancetrack was also used.
For KITTI, initial training involved 80 epochs, followed by 25 in the second stage. The third stage consisted of 30 epochs with TBSP or 70 without. Additionally, approximately 5k images from BDD100k [48] were selected for additional training.
4.3 Main Result
The comprehensive benchmarking in Table 1 demonstrates DecoderTracker’s dual superiority in training efficiency and inference speed over MOTR-series methods. This design achieves 68.7 training hours on a single GPU, equivalent to 7.1× higher GPU-hour efficiency versus MOTR’s 768 GPU-hour consumption (8 GPUs × 96 hours). The phased training regimen—12.1 hours for feature embedding, 30.8 hours for temporal modeling, and 43.4 hours for TBSP-enhanced prediction—systematically constructs tracking capabilities while maintaining 83% shorter wall-clock time than competitors. Inference performance reveals more pronounced advantages: DecoderTracker+ attains 28.8 FPS, 4.1× faster than MOTRv2’s 6.9 FPS,
For DanceTrack ,the experimental results shown in Table 2 reveal that DecoderTracker and its enhanced variant DecoderTracker+ achieve comparable precision levels across core evaluation metrics. Both models demonstrate superior detection accuracy over existing end-to-end methods, with DetA scores of 78.7 and 79.0 respectively, significantly outperforming MOTR’s 73.5. Their MOTA metrics (87.8 and 88.1) approach SOTA non-end-to-end methods like ByteTrack (89.5), while maintaining balanced HOTA performance at 54.2 and 54.5. This parity suggests that architectural refinements between the two variants primarily yield marginal improvements rather than fundamental performance shifts. Notably, both configurations exhibit identical IDF1 scores (54.2 and 54.4), confirming consistent identity preservation capabilities. The minimal performance delta between variants implies optimized parameter utilization rather than architectural overhauls. When extended with supplementary training data (DecoderTracker* variants), both models achieve substantial gains, attaining 91.1-91.9 MOTA and 60.8-61.2 HOTA. While their AssA scores (35.6-35.7) remain below MOTR’s 40.2, the detection-centric advantages manifest through DetA superiority over all end-to-end and most non-end-to-end competitors. These findings position the DecoderTracker framework as a robust baseline for end-to-end tracking, achieving detection reliability comparable to sophisticated multi-stage systems while preserving model efficiency.